A study of the time spent shopping in a supermarket for a market basket of 20 specific items showed an approximately uniform distribution between 20 minutes and 40 minutes.

  1. What is the probability that the shopping time will be between 27 and 45 minutes?
    1. 0.35
    2. 0.65
    3. 0.90
    4. 0.15
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

there are 20 items,

the uniform distribution is between 20 and 40 minutes

P(27 and 45) = ?

P(27<X<45) = (x-a)/(b-a), where b and a are 40 and 20. Please notice that the limit of 40 will be applied for values over 40. So, if we will try to find out the the probability of range 27 to 40 instead of 27 to 45, as 40 is our upper range.

So, P(27<X<40) = (40-27)/(40-20) = 13/20 = .65

Answer is B. 0.65